### v1

@@
Diffusion Large Language Models (dLLMs) break the rigid left-to-right constraint of traditional LLMs, enabling token generation in arbitrary orders. Intuitively, this flexibility implies a solution space that strictly supersets the fixed autoregressive trajectory, theoretically unlocking superior reasoning potential for general tasks like mathematics and coding. Consequently, numerous works have leveraged reinforcement learning (RL) [^1] In this paper, unless otherwise specified, we use the term RL to refer specifically to Reinforcement Learning with Verifiable Rewards (RLVR), as it is the predominant paradigm for enhancing the reasoning capabilities of dLLMs. to elicit the reasoning capability of dLLMs. In this paper, we reveal a counter-intuitive reality: arbitrary order generation, in its current form, narrows rather than expands the reasoning boundary of dLLMs. We find that dLLMs tend to exploit this order flexibility to bypass high-uncertainty tokens that are crucial for exploration, leading to a premature collapse of the solution space. This observation challenges the premise of existing RL approaches for dLLMs, where considerable complexities, such as handling combinatorial trajectories and intractable likelihoods, are often devoted to preserving this flexibility. We demonstrate that effective reasoning is better elicited by intentionally forgoing arbitrary order and applying standard Group Relative Policy Optimization (GRPO) instead. Our approach, JustGRPO, is minimalist yet surprisingly effective (e.g., 89.1% accuracy on GSM8K) while fully retaining the parallel decoding ability of dLLMs. Project page: https://nzl-thu.github.io/the-flexibility-trap
Figure 1: Less flexibility unlocks better reasoning potential. Left: We observe a counter-intuitive phenomenon where restricting dLLMs to standard Autoregressive (AR) order expands the reasoning solution space. Right: Motivated by this, we propose “JustGRPO”. By foregoing complex arbitrary-order adaptations and adopting standard GRPO, we effectively elicit the reasoning capability of dLLMs.
@@ 1 Introduction
While the efficiency gains of parallel decoding have been well-established Wu et al. (2025b); Labs et al. (2025); DeepMind (2025); Song et al. (2025); Wu et al. (2025a), the implications of arbitrary-order generation remain less explored Ye et al. (2024); Kim et al. (2025). Theoretically, the unconstrained generation order constitutes a superset of the fixed autoregressive trajectory. This flexibility naturally suggests a potential for superior reasoning: in general reasoning tasks like mathematics and coding, such freedom could unlock non-sequential problem-solving paths inaccessible to standard left-to-right models. As a result, recent works have increasingly adopted RL to elicit reasoning capabilities of dLLMs Zhao et al. (2025); Gong et al. (2025); Wang et al. (2025a); Ou et al. (2025).
@@
We attribute this counter-intuitive phenomenon to the way model handles uncertainty. Reasoning is inherently non-uniform: it typically hinges on sparse “forking tokens”, i.e., connectives like “Therefore” or “Since” which do not merely continue a sentence but fundamentally steer the logical trajectory into distinct branches Wang et al. (2025c); Cheng et al. (2025); Huang et al. (2025a). At these forks, the reasoning path diverges, naturally manifesting as localized spikes in entropy Wang et al. (2025c). Standard AR decoding compels the model to confront this uncertainty (Figure 2a). By sampling exactly at the fork, the model is able to explore different reasoning paths, thereby preserving the diversity of the generated rationales. Arbitrary-order generation, however, allows the model to bypass these hard decisions (Figure 2b). It prioritizes low-entropy completions first. By the time the model returns to infill the bypassed forks, the established bidirectional context has already severely constrained the potential branches. The ambiguity is prematurely resolved, and the original high entropy is suppressed. We term this phenomenon entropy degradation. Effectively, the model trades the exploration of diverse reasoning paths for the greedy optimization of local consistency.
@@
To this end, we propose a return to simplicity with Just GRPO. We demonstrate that eliciting reasoning potential for general reasoning tasks does not require complex, diffusion-specific RL adaptations. Instead, it is best achieved by simply treating the dLLM as an AR model during RL training. This allows us to apply standard GRPO Shao et al. (2024) without bells and whistles, turning an otherwise intractable optimization with unstable approximations into a well-defined task.
JustGRPO is surprisingly effective. On complex reasoning benchmarks, it achieves competitive results (e.g., 89.1% accuracy on GSM8K, 45.1% on MATH), surpassing methods that rely on complex diffusion-specific RL. Crucially, while we train with AR constraints to maximize reasoning potential, the model retains dLLMs’ ability of efficient parallel decoding at inference time. By returning to basic left-to-right ordering, this work encourages a re-evaluation of arbitrary versus AR order in the development of next-generation language models.
@@ 2.1 Diffusion Large Language Models
$q(x_{t}^{k}\mid x_{0}^{k})=\begin{cases}\texttt{[MASK]},&\text{with prob }t,\\ x_{0}^{k},&\text{with prob }1-t.\end{cases}$
Unlike Gaussian diffusion, MDMs directly predict the clean token. A neural network $p_{\theta}(x_{0}\mid x_{t})$ estimates the original token distribution at masked positions. During inference, generation starts from $x_{1}$ (all [MASK]) and iteratively unmasks a subset of tokens based on heuristics such as confidence scores, updating $x_{t}\rightarrow x_{t-\Delta t}$ until completion. As a special case, autoregressive generation can be recovered by always unmasking the leftmost token. The model is trained by minimizing the Negative Evidence Lower Bound, which reduces to a weighted cross entropy loss over masked tokens:
$\mathcal{L}_{\mathrm{MDM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{t}\sim q(x_{t}\mid x_{0})}\left[\frac{1}{t}\sum_{k=1}^{L}\mathbf{1}[x_{t}^{k}=\texttt{[MASK]}]\,\log p_{\theta}(x_{0}^{k}\mid x_{t})\right].$
@@ 2.2 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) is a reinforcement learning algorithm that avoids value function estimation by using group level reward normalization. It is typically applied to autoregressive policies $\pi_{\theta}(o_{k}\mid o_{<k},q)$. For each query $q$, GRPO samples a group of $G$ outputs $\{o_{1},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. An advantage $A_{i}$ is computed by standardizing the reward $r(o_{i})$ against the group statistics: $A_{i}=(r(o_{i})-\mu_{G})/\sigma_{G}$, where $\mu_{G}$ and $\sigma_{G}$ are the group mean and standard deviation. The GRPO objective maximizes a clipped surrogate function with a KL regularization term:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{L_{i}}\sum_{k=1}^{L_{i}}\left(\min\left(\rho_{i,k}A_{i},\text{clip}(\rho_{i,k},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right)\right]$ (1)
@@ 2.3 Pass@$k$ as a Proxy for Reasoning Potential
Accordingly, Pass@$k$ has been widely established as a standard measure of a model’s reasoning potential Yue et al. (2025); Liu et al. (2025); Zhang et al. (2025). It measures the probability that at least one correct solution is generated within $k$ independent samples, effectively delineating the upper bound of the solution space accessible to the RL optimizer. Formally, following the unbiased estimator formulation Chen (2021); Yue et al. (2025), given $n$ samples where $c$ are correct, Pass@$k$ is calculated as:
@@
Figure 3: Reasoning potential measured by Pass@$k$. While arbitrary order is competitive in single-shot settings ($k=1$), it exhibits notably flatter scaling curves compared to AR Order.
@@ 3 The Flexibility Trap
In this section, we rigorously test whether the theoretical flexibility of arbitrary-order generation translates into higher reasoning potential. We compare two decoding modes: Arbitrary Order, which follows standard diffusion decoding with low-confidence remasking Nie et al. (2025); Zhao et al. (2025); Wu et al. (2025b), and AR Order, where arbitrary-order flexibility is disabled and generation is strictly constrained to left-to-right decoding. We adopt the commonly used experimental setup in prior diffusion LLM work: a maximum of 256 tokens decoded over 256 steps, a semi-autoregressive block size of 32. Sampling temperature is set to 0.6 as in Yue et al. (2025). The prompt template follows Zhao et al. (2025). Results under alternative temperatures and sampling configurations are deferred to Appendix B.
@@ 3.1 Arbitrary Order Limits Reasoning Potential
We first evaluate the reasoning potential using the Pass@$k$ metric on three representative dLLMs: LLaDA-Instruct Nie et al. (2025), Dream-Instruct Ye et al. (2025), and LLaDA 1.5 Zhu et al. (2025) on four reasoning benchmarks: GSM8K, MATH500, HumanEval, and MBPP. As shown in Figure 3, while arbitrary order often achieves competitive performance at $k=1$, it exhibits a notably flatter scaling curve compared to the AR mode. As $k$ increases, the AR mode demonstrates a stronger capacity to uncover correct solutions.
Figure 4: Solution space coverage measured by Pass@$1024$. The reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR Order.
@@
One might hypothesize that arbitrary order explores a different solution space, albeit less efficiently, which could account for its lower reasoning potential. We test this by analyzing solution coverage at $k=1024$ using LLaDA-Instruct. Figure 4 presents a stark reality: the reasoning traces generated by arbitrary order are largely a strict subset of those generated by AR. On HumanEval, AR solves 21.3% of the problems that arbitrary order misses, whereas the reverse is negligible (0.6%). This indicates that the flexible decoding process rarely unlocks genuinely new solutions. Instead, it appears to retreat into a more conservative subset of the AR solution space.
@@ 3.2 Mechanism: The Entropy Degradation
To understand why the theoretically superior solution space of dLLMs collapses in practice, we examine more closely how the two decoding modes handle uncertainty. In AR order, the model is constrained to strictly resolve the left-most unknown token at each step, forcing the model to confront uncertainty as it arises. In contrast, arbitrary order adaptively selects tokens to update based on model confidence, preferentially generating “easy” tokens with high certainty while bypassing “hard” ones. Inspecting the frequently bypassed tokens reveals a clear pattern: As shown in Figure 5, the diffusion sampler disproportionately defers logical connectives and transition markers such as “Therefore”, “Thus”, and “Since”. Prior work has shown that such tokens often have high entropy (which also holds true in dLLMs; see Figure 6), and act as “reasoning sparks” or “logical forks”, functioning as branching points that determine subsequent reasoning directions Wang et al. (2025c); Cheng et al. (2025); Wang et al. (2025b); Huang et al. (2025a). In conventional language models, keeping these tokens in high-entropy state is critical for effective exploration of the reasoning space Wang et al. (2025c).
Figure 6: Entropy degradation. While the global average entropy of Arbitrary Order remains comparable to AR (dashed lines), the entropy at logical forks drops significantly (blue bars).
@@
This adaptive behavior comes at a cost: the premature collapse of reasoning possibilities. We measure the entropy of these pivotal connectors at decoding. In AR order, these tokens maintain high entropy, reflecting a genuine branching point where multiple logical paths remain viable. In contrast, arbitrary order exhibits a sharp decrease in entropy (Figure 6, with more results in Appendix B). By deferring the logical connector, the model commits to a specific future outcome that is generated based on its inherent inductive biases before deciding the logic that leads there. When the model eventually returns to fill in the bypassed connector, it is no longer making a navigational decision at a fork; it is more like selecting the connector that grammatically bridges the gap to its pre-generated conclusion. The decoding process thus implicitly shifts from reasoning exploration into semantic pattern matching. We term this phenomenon entropy degradation.
@@ Conclusion.
In summary, the flexibility of arbitrary order serves as a mechanism for inference-time exploitation rather than reasoning exploration. By bypassing high-uncertainty tokens, the model effectively collapses the solution space to a safe, low-entropy path, squeezing out slightly better single-shot coherence at the expense of reasoning potential. Autoregressive models, by contrast, lack this bypassing capability and are therefore forced to sample directly from high-entropy distributions at logical forks. It is precisely this inability to circumvent critical decision points that prevents the premature narrowing of the search space and preserves the reasoning potential.
@@ 4.1 The Flexibility Tax in dLLMs’ RL
Existing diffusion RL methods operate under the premise that the policy must optimize over the full combinatorial space of denoising trajectories $\mathcal{T}$ to preserve the flexibility of arbitrary order generation. This design choice, while conceptually general, introduces several fundamental challenges.
@@
Autoregressive models factorize the sequence likelihood as $\log\pi(o)=\sum_{t}\log\pi(o_{t}\mid o_{<t},q)$, whereas dLLMs require marginalization over all valid denoising trajectories, $\pi_{\theta}(o\mid q)=\sum_{\tau\in\mathcal{T}}\pi_{\theta}(o,\tau\mid q)$. For a sequence of length $N$, the trajectory space grows as $|\mathcal{T}|=O(N!)$, rendering exact likelihood computation infeasible and forcing existing methods to rely on ELBO-based surrogates rather than the true objective Wang et al. (2025a); Ou et al. (2025).
@@
Even with an accurate likelihood approximation, a more subtle issue persists. In practice, rollout samples are produced by heuristic-guided policies $o\sim\pi_{\theta}^{\text{heur}}(o\mid q)$ to explore the combinatorial space. However, the ELBO objective still targets the likelihood of the original model distribution $\pi_{\theta}(o\mid q)$, rather than that of the heuristic-guided sampler $\pi_{\theta}^{\text{heur}}(o\mid q)$, leading to a critical mismatch between sampling and optimization that can degrade performance Schulman et al. (2015).

### v2

@@
Diffusion Large Language Models (dLLMs) break the rigid left-to-right constraint of traditional LLMs, enabling token generation in arbitrary orders. Intuitively, this flexibility implies a solution space that strictly supersets the fixed autoregressive trajectory, theoretically unlocking superior reasoning potential for general tasks like mathematics and coding. Consequently, numerous works have leveraged reinforcement learning (RL) [^1] In this paper, unless otherwise specified, we use the term RL to refer specifically to Reinforcement Learning with Verifiable Rewards (RLVR), as it is the predominant paradigm for enhancing the reasoning capabilities of dLLMs. to elicit the reasoning capability of dLLMs. In this paper, we reveal a counter-intuitive reality: arbitrary order generation, in its current form, narrows rather than expands the reasoning boundary of dLLMs. We find that dLLMs tend to exploit this order flexibility to bypass high-uncertainty tokens that are crucial for exploration, leading to a premature collapse of the solution space. This observation motivates a rethink of RL approaches for dLLMs, where considerable complexities, such as handling combinatorial trajectories and intractable likelihoods, are often devoted to preserving this flexibility. We demonstrate that effective reasoning can be better elicited by intentionally forgoing arbitrary order and applying standard Group Relative Policy Optimization (GRPO) instead. Our approach, JustGRPO, is minimalist yet surprisingly effective (e.g., 89.1% accuracy on GSM8K) while fully retaining the parallel decoding ability of dLLMs. Project page: https://nzl-thu.github.io/the-flexibility-trap
Figure 1: Less flexibility unlocks better reasoning potential. Left: We observe a counter-intuitive phenomenon where restricting dLLMs to standard Autoregressive (AR) order expands the reasoning solution space. Right: Motivated by this, we propose “JustGRPO”. By foregoing complex arbitrary-order adaptations and adopting just GRPO, we elicit the reasoning capability of dLLMs more effectively. Results reported with generation length of 256, more results in Table 1.
@@ 1 Introduction
While the efficiency gains of parallel decoding have been well-established Wu et al. (2025b); Labs et al. (2025); DeepMind (2025); Song et al. (2025); Wu et al. (2025a), the implications of arbitrary-order generation remain less explored Ye et al. (2024); Kim et al. (2025). In theory, the unconstrained generation order constitutes a superset of the fixed autoregressive trajectory, unlocking non-sequential problem-solving paths. Ideally, such advantage could facilitate more advanced problem-solving abilities that benefit general reasoning tasks like mathematics and coding. As a result, recent works have increasingly adopted RL to elicit reasoning capabilities of dLLMs Zhao et al. (2025); Gong et al. (2025); Wang et al. (2025a).
@@
We attribute this counter-intuitive phenomenon to the way model handles uncertainty. Reasoning is inherently non-uniform: it typically hinges on sparse “forking tokens”, i.e., connectives like “Therefore” or “Since” which do not merely continue a sentence but fundamentally steer the logical trajectory into distinct branches Wang et al. (2025c); Cheng et al. (2025); Huang et al. (2025a). At these forks, the reasoning path diverges, naturally manifesting as localized spikes in entropy Wang et al. (2025c). Standard AR decoding compels the model to confront this uncertainty (Figure 2a). By sampling exactly at the fork, the model is able to explore different reasoning paths, thereby preserving the diversity of the generated rationales. Arbitrary-order generation, however, allows the model to bypass these hard decisions (Figure 2b). It prioritizes low-uncertainty completions first. By the time the model returns to infill the bypassed forks, the established future context has already severely constrained the potential branches. The ambiguity is prematurely resolved, and the original high entropy is suppressed. We term this phenomenon entropy degradation. Effectively, the model trades the exploration of diverse reasoning paths for the greedy optimization of local consistency.
@@
To this end, we propose a return to simplicity with JustGRPO. We demonstrate that eliciting reasoning potential for general reasoning tasks does not require complex, diffusion-specific RL adaptations. Instead, it can be better achieved by simply treating the dLLM as an AR model during RL training. This allows us to apply standard GRPO Shao et al. (2024) without bells and whistles, turning an otherwise intractable optimization with unstable approximations into a well-defined task.
JustGRPO is surprisingly effective. On complex reasoning benchmarks, it achieves competitive results (e.g., 89.1% accuracy on GSM8K, 45.1% on MATH-500), surpassing methods that rely on complex diffusion-specific RL. Crucially, while we train with AR constraints to maximize reasoning potential, the model retains dLLMs’ ability of efficient parallel decoding at inference time. By returning to basic left-to-right ordering, this work encourages a re-evaluation of arbitrary versus AR order in the development of next-generation language models.
@@ 2.1 Diffusion Large Language Models
$q(x_{t,k}\mid x_{0,k})=\begin{cases}\texttt{[MASK]},&\text{with prob }t,\\ x_{0,k},&\text{with prob }1-t,\end{cases}$ (1)
where $k$ is the token index. A neural network $p_{\theta}(x_{0}\mid x_{t})$ estimates the original token distribution at masked positions. During inference, generation starts from $x_{1}$ (all [MASK]) and iteratively unmasks a subset of tokens based on heuristics such as confidence scores, updating $x_{t}\rightarrow x_{t-\Delta t}$ until completion. As a special case, dLLMs can also perform generation autoregressively by always unmasking the leftmost token. The model is trained by minimizing the Negative Evidence Lower Bound, which reduces to a weighted cross entropy loss over masked tokens:
$\mathcal{L}_{\mathrm{MDM}}(\theta)=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{t}\sim q(x_{t}\mid x_{0})}\left[\frac{1}{t}\sum_{k=1}^{L}\mathbf{1}[x_{t,k}=\texttt{[MASK]}]\,\log p_{\theta}(x_{0,k}\mid x_{t})\right].$ (2)
@@ 2.2 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) is a reinforcement learning algorithm that avoids value function estimation by using group level reward normalization. It is typically applied to autoregressive policies $\pi_{\theta}(o_{k}\mid o_{<k},q)$. For each query $q$, GRPO samples a group of $G$ outputs $\{o_{1},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. An advantage $\hat{A}_{i,k}$ is computed by standardizing the reward $r(o_{i})$ against the group statistics: $\hat{A}_{i,k}=(r(o_{i})-\mu_{G})/\sigma_{G}$, where $\mu_{G}$ and $\sigma_{G}$ are the group mean and standard deviation. The GRPO objective maximizes a clipped surrogate function with a KL regularization term:
$\mathcal{J}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\left(\min\left(\rho_{i,k}\hat{A}_{i,k},\text{clip}(\rho_{i,k},1-\epsilon,1+\epsilon)\hat{A}_{i,k}\right)-\beta D_{\text{KL}}\right)\right]$ (3)
@@ 2.3 Pass@$k$ as a Proxy for Reasoning Potential
Accordingly, Pass@$k$ has been widely established as a standard measure of a model’s reasoning potential Yue et al. (2025); Liu et al. (2025); Zhang et al. (2025). It measures the probability that at least one correct solution is generated within $k$ independent samples, effectively representing the upper bound of the solution space accessible to RL optimization. Formally, following the unbiased estimator formulation Chen (2021); Yue et al. (2025), given $n$ samples where $c$ are correct, Pass@$k$ is calculated as:
@@
Figure 3: Reasoning potential measured by Pass@$k$. While arbitrary order is competitive in single-shot settings ($k=1$), it exhibits flatter scaling curves compared to AR Order.
@@ 3 The Flexibility Trap
In this section, we rigorously test whether the flexibility of arbitrary-order generation translates into higher reasoning potential. We compare two decoding modes: Arbitrary Order, which follows standard diffusion decoding with low-confidence remasking Nie et al. (2025); Zhao et al. (2025); Wu et al. (2025b), and AR Order, where arbitrary-order flexibility is disabled and generation is strictly constrained to left-to-right decoding. We adopt the commonly used experimental setup in prior diffusion LLM work: a maximum of 256 tokens decoded over 256 steps, a semi-autoregressive block size of 32. Sampling temperature is set to 0.6 as in Yue et al. (2025). The prompt template follows Zhao et al. (2025). Results under alternative temperatures and sampling configurations are deferred to Appendix B.
@@ 3.1 Arbitrary Order Limits Reasoning Potential
We first evaluate the reasoning potential using the Pass@$k$ metric on three representative dLLMs: LLaDA-Instruct Nie et al. (2025), Dream-Instruct Ye et al. (2025), and LLaDA 1.5 Zhu et al. (2025) on four reasoning benchmarks: GSM8K, MATH-500, HumanEval, and MBPP. As shown in Figure 3, while arbitrary order often achieves competitive performance at $k=1$, it exhibits a notably flatter scaling curve compared to the AR mode. As $k$ increases, the AR mode demonstrates a stronger capacity to uncover correct solutions.
Figure 4: Solution space coverage measured by Pass@$1024$. The problems solvable by arbitrary order (AO) are largely a subset of those solved by AR Order.
@@
One might wonder whether arbitrary order explores a different solution space, albeit less efficiently, which could account for its lower reasoning potential. We test this by analyzing solution coverage at $k=1024$ using LLaDA-Instruct. Figure 4 shows that the problems solvable by arbitrary order are largely a subset of those solved by AR. On HumanEval, 21.3% of problems are solved exclusively by AR, whereas the reverse is negligible (0.6%). This indicates that the flexible decoding process rarely unlocks genuinely new solutions. Instead, it appears to retreat into a more conservative subset of the AR solution space.
@@ 3.2 Mechanism: The Entropy Degradation
To understand why the theoretically superior solution space of dLLMs collapses in practice, we examine more closely how the two decoding modes behave. In AR order, the model is constrained to strictly resolve the left-most unknown token at each step, forcing the model to confront uncertainty as it arises. In contrast, arbitrary order adaptively selects tokens to update based on model confidence, preferentially generating “easy” tokens with high certainty while bypassing “hard” ones. Inspecting the frequently bypassed tokens reveals a clear pattern: As shown in Figure 5, the diffusion sampler disproportionately defers logical connectives and transition markers such as “Therefore”, “Thus”, and “Since”. Prior work has shown that such tokens often have high entropy (which also holds true in dLLMs; see Figure 6), and act as “reasoning sparks” or “logical forks”, functioning as branching points that determine subsequent reasoning directions Wang et al. (2025c); Cheng et al. (2025); Wang et al. (2025b); Huang et al. (2025a). In conventional language models, keeping these tokens in high-entropy state is critical for effective exploration of the reasoning space Wang et al. (2025c).
Figure 6: Entropy degradation. While the global average token entropy (dashed lines) of arbitrary order remains comparable to AR, the entropy at logical forks (blue bars) drops significantly.
@@
How does the adaptive behavior of arbitrary order affect these logical forking tokens? We measure the entropy of them when they are decoded in Figure 6 (more results in Appendix C). In AR order, these tokens generally maintain high entropy when decoded, indicating a relatively fruitful branching point where multiple reasoning paths remain viable Wang et al. (2025c); Cheng et al. (2025). In contrast, arbitrary order exhibits a sharp decrease in entropy. By deferring the logical connectors, the model prioritizes generating easier future tokens before deciding the logic connections that lead to them. When the model eventually returns to fill in the bypassed connectors, the navigational freedom is notably constrained; the process acts less as an open-ended navigational decision at a fork, and more as a retrospective alignment to bridge the gap to the pre-generated conclusion. We term this phenomenon entropy degradation.
@@ Conclusion.
In summary, the flexibility of arbitrary order serves as a mechanism for inference-time exploitation rather than reasoning exploration. By bypassing high-uncertainty tokens, the model effectively collapses the solution space to a safer, lower-entropy path, squeezing out slightly better single-shot coherence at the expense of reasoning potential. Autoregressive ordering, by contrast, lack this bypassing capability and are therefore forced to sample directly from high-entropy distributions at logical forks. It is precisely this inability to circumvent critical decision points that prevents the premature narrowing of the search space and preserves the reasoning potential.
@@ 4.1 The Flexibility Tax in dLLMs’ RL
Existing diffusion RL methods operate under the premise that the policy must accommodate the combinatorial space of denoising trajectories $\mathcal{T}$ to preserve arbitrary order generation. This design choice, while conceptually general, introduces several fundamental challenges:
@@
Autoregressive models factorize the sequence likelihood as $\log\pi(o)=\sum_{t}\log\pi(o_{t}\mid o_{<t},q)$, whereas dLLMs require marginalization over all valid denoising trajectories, $\pi_{\theta}(o\mid q)=\sum_{\tau\in\mathcal{T}}\pi_{\theta}(o,\tau\mid q)$. For a sequence of length $N$, the trajectory space grows as $|\mathcal{T}|=O(N!)$, rendering exact likelihood computation infeasible and forcing existing methods to rely on ELBO-based surrogates rather than the original objective Wang et al. (2025a); Ou et al. (2025).
@@
Even with an accurate likelihood approximation, a more subtle issue persists. In practice, rollout samples are typically produced by confidence-based generation $o\sim\pi_{\theta}^{\text{conf}}(o\mid q)$ to navigate the combinatorial space. However, the ELBO objective still targets the likelihood of the original model distribution $\pi_{\theta}(o\mid q)$, rather than that of the actual sampling policy $\pi_{\theta}^{\text{conf}}(o\mid q)$, leading to a mismatch between rollout and optimization that can degrade performance Schulman et al. (2015).